# Fermion-qubit fault-tolerant quantum computing

### v1

@@
Fermion-qubit fault-tolerant quantum computing
@@
(November 13, 2024)
Simulating the dynamics of electrons and other fermionic particles in quantum chemistry [1], material science [2], and high-energy physics [3] is one of the most promising applications of fault-tolerant quantum computers. However, the overhead in mapping time evolution under fermionic Hamiltonians to qubit gates [4, 5, 6, 7] renders this endeavor challenging [8, 9]. We introduce fermion-qubit fault-tolerant quantum computing, a framework which removes this overhead altogether. Using native fermionic operations we first construct a repetition code which corrects phase errors only. We then engineer a fermionic color code which corrects for both phase and loss errors. We show how to realize a universal fermionic gate set in this code, including transversal Clifford gates. Interfacing with qubit color codes we realize qubit-fermion fault-tolerant computation, which allows for qubit-controlled fermionic time evolution, a crucial subroutine in state-of-the-art quantum algorithms for simulating fermions [10, 11, 12, 13]. We show how our framework can be implemented in neutral atoms, overcoming the apparent inability of neutral atoms to implement non-number-conserving gates by introducing a neutral-atom braiding gate using photodissociation of bosonic molecules. As an application, we consider the fermionic fast Fourier transform, an important subroutine for simulating crystalline materials [12], finding an exponential improvement in circuit depth from $\mathcal{O}(N)$ to $\mathcal{O}(\log(N))$ with respect to lattice site number $N$ and a linear improvement from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log(N))$ in Clifford gate complexity compared to state-of-the-art qubit-only approaches [14]. Our work opens the door to fermion-qubit fault-tolerant quantum computation in platforms with native fermions such as neutral atoms, quantum dots and donors in silicon, with applications in quantum chemistry, material science, and high-energy physics.
@@
Solving the dynamics of many strongly-interacting fermionic quantum particles is central to numerous scientific and technological challenges, including explaining the dynamics of quarks in the early universe, understanding the reaction dynamics of molecules, and elucidating high-temperature superconductivity. While qubit quantum computers promise to help solve these open problems, simulating the evolution of fermions encoded in qubits encounters a large overhead in both the number of two-qubit gates and the circuit depth [5, 6, 7], which means that a solution of fermionic problems with near-term [8, 9] and even fault-tolerant quantum computers [12] is extremely challenging. It is possible to circumvent this issue by using quantum simulation platforms which host controllable fermionic degrees of freedom, such as quantum dots [15], donors in silicon [16], and light-trapped neutral atoms [17, 18, 19, 20]. Using those platforms, analog simulations of condensed-matter models have been achieved [21, 22, 23, 24, 25] and proposals for studying quantum chemistry in an analog [26] or digital [27] way have been made. However, turning these fermionic platforms from a tool of scientific exploration into a computational tool requires error-correction and fault-tolerance techniques, for which no practical scheme has been found so far.
Here, we close this gap by introducing a framework for fermion-to-fermion fault-tolerant quantum computation. We first prove a theorem that fermionic error correction is not possible using only number-conserving gates, which are most easily implemented in experimental platforms. We then show how to circumvent this theorem by using a non-number-conserving gate and concepts from Majorana error correction. This enables us to introduce error-correcting codes for which we explicitly show how to perform a universal gate set including fault-tolerant transversal gates. To enable usage of state-of-the-art quantum algorithms requiring qubit-controlled gates [10], we introduce qubit-fermion fault-tolerant computation. We apply our framework to time evolution under the Hamiltonian of a crystalline material and study the advantages of our framework compared to all-qubit quantum computation. Finally, we show how to realize these codes and their operations in neutral atoms by removing the previously assumed restriction to number-conserving gates [27] in neutral atoms and introducing the physics behind the implementation of a non-number-conserving braiding gate.
@@
The proof (Methods) relies on the fact that codestates that are eigenstates of the total fermion number operator need to be eigenstates with the same number eigenvalue due to the Knill-Laflamme conditions. This means that logical operators cannot change the total fermion number and therefore need to be even weight in fermionic creation/annihilation operators. Hence, logical operators are bosonic.
@@
For implementing active error correction, we need to measure where errors happened and correct for them, which we show how to do in Fig. 2b. We use our qubit-fermion gate to map a measurement of the fermion density $n$ to a measurement of the phase of the qubit, which we transform to a measurement of the population of the qubit using Hadamard gates. To measure the stabilizer instead of $n$ on the fermions we conjugate the qubit-fermion gate with two braiding gates as well as phase gates. An ancilla qubit is used for each of the stabilizers. Dependent on the measurement outcome, $\mathrm{Z}^{\mathrm{f}}$  gates are performed on the fermions to correct for the errors. This scheme shows the first advantage of using qubits: destructive qubit readout is faster than a non-destructive $n$ measurement [35]. To prepare a logical codestate |$\ket{0}_{\mathrm{L}}$, we use a chain of braiding gates—the $\mathrm{S}^{\dagger}$ gates effectively braid the two Majoranas on the same physical site, Fig. 2c. Note that we do not error correct the qubit ancillas, which is reasonable because their dephasing time is much longer than that of the fermions and we only act with a handful of gates on them before measuring.
@@
We now show how to construct an error-correcting code correcting for both phase errors $E=n$ and atom-loss errors $E=p$ and show how to perform fault-tolerant gadgets. In the context of encoding qubits in Majorana nanowire platforms, Bravyi, Terhal, and Leemhuis [31] noted that a particular class of qubit error correcting codes, so-called weakly self-dual Calderbank-Shor-Steane (CSS) codes, can be converted into a Majorana stabilizer code by replacing qubit operators in the stabilizers by Majorana fermions. Motivated by the prospect of high error-resilience for codes with odd-weight logical operators, they noted that for codes with an odd number of physical Majorana fermions, a single odd-weight logical operator can be encoded. This operator is given by the total parity, i.e. the product of all Majorana operators. Our key observation is that, because there is an odd number of physical Majorana fermions, this logical operator is in fact a logical Majorana fermion operator. Even more importantly, our framework using complex physical fermions enables the encoding of two such Majorana codes in the same set of physical fermions. It turns out that a general prescription to constructing complex fermion codes from weakly self-dual CSS codes is to replace the $Z$ operators in the qubit code stabilizers by $\gamma$ operators and the qubit $X$ operators by $\tilde{\gamma}$ operators, see Fig. 3a. The logical Majorana operators are then given by the two parities $\gamma^{\mathrm{L}}=\prod_{i=1}^{N}\gamma_{i}$ and $\tilde{\gamma}^{\mathrm{L}}=\prod_{i=1}^{N}\tilde{\gamma}_{i}$, from which we construct a single logical fermion operator $c^{\dagger}=\frac{1}{2}\left(\gamma^{\mathrm{L}}+i\tilde{\gamma}^{\mathrm{L}}\right)$. As an example, we consider the family of triangular honeycomb Majorana color codes, which using our procedure above, encodes a single logical fermion with odd distance $d$ into $N=(3d^{2}+1)/4$ physical fermions. In particular, $d=3$ is the fermionic version of the Steane code [36], Fig. 3a. Such codes can correct $d-2$ loss or phase errors.
@@
Figure 3: Fermion-qubit fault-tolerant quantum computation. a, Construction of the fermionic Steane code, the smallest member of the family of color codes we consider. Stabilizers are obtained by replacing products of $X$ ($Z$) operators in the qubit code with products of $\gamma$ ($\tilde{\gamma}$), respectively. Corners on areas with the same color are part of a stabilizer as shown for the bottom left area. b, Mid-circuit measurement of $-\gamma_{1}\gamma_{2}\gamma_{3}\gamma_{4}$ (counting sites from bottom to top). By inverting the $\mathrm{BRAID}$s (i.e. $\mathrm{BRAID}_{12}\rightarrow\mathrm{BRAID}_{21}$), we measure $-\tilde{\gamma}_{1}\tilde{\gamma}_{2}\tilde{\gamma}_{3}\tilde{\gamma}_{4}$, the other stabilizer. c, Neutral-atom implementation of transversal $\mathrm{CZ}^{\mathrm{qf}}$  gates, performed by interleaving a logical qubit and fermion and applying a global Rydberg laser. This implements the circuit shown in subfigure e. d, Transversal $\mathrm{CZ}^{\mathrm{f}}$. This is implemented by the logical qubit in subfigure c with a logical fermion. e, Transversal qubit-fermion $\mathrm{CZ}^{\mathrm{qf}}$. Transversal f, $\mathrm{BRAID}$  and g, $\mathrm{S}^{\mathrm{f}}$  gates are similarly implemented by parallel operations. h, $\mathrm{T}^{\mathrm{f}}$gate. Implemented by using a T gate applied on an ancilla qubit, which is then mapped onto the fermion. All the operations shown here are straightforwardly generalized to codes with larger (odd) $d$ (see Methods).
@@
Magically, this fermion Steane code yields highly intuitive and easy-to-implement fault-tolerant logical gates. Of central utility is the fact that we can use qubits for controlled operations. We use this fact to measure stabilizers using the Hadamard test, as shown in Fig. 3b. From these stabilizers, we can reconstruct whether a $\tilde{\gamma}$ or $\gamma$ error (or both) happened on a fermion. Such single-Majorana errors are corrected by coupling an ancilla site $a$ prepared in |$\ket{0}$ to the site with the error, acting with $(\mathrm{BRAID}_{ai})^{2}=\gamma_{i}\tilde{\gamma}_{a}$ and then measuring the ancilla. The measurement circuit in Fig. 3b is made fault-tolerant by for example using the Shor scheme, i.e. replacing the single ancillary qubit with a set of qubits in a cat state and repeatedly measuring the stabilizers. This stabilizer measurement is also used for fault-tolerant preparation of the logical codestates from a trivial starting state. Note that the correction of stabilizers is optional for both state preparation and error correction—instead of enforcing stabilizers to be +1, their value can instead be tracked [37].
@@
Transversal logical gates are implemented, similar to qubit Steane codes, by applying the corresponding physical gate in parallel on all atoms, c.f. Fig. 3c-g (Methods). This set of gates turns out to be the Majorana Clifford gates $\mathrm{BRAID}$, $\mathrm{CZ}^{\mathrm{f}}$, $\mathrm{S}^{\mathrm{f}}$, which in analogy to qubit Clifford gates map a single Majorana operator string to another single Majorana string. These transversal Clifford gates are inherently fault-tolerant as an error on one physical codeblock can only spread to one physical fermion in the other codeblock.
@@
One of the key tools enabled by our approach is the interfacing of our fermionic Steane code with a qubit Steane code (with matching $N$) using the qubit-fermion $\mathrm{CZ}^{\mathrm{qf}}$ gate: by applying physical $\mathrm{CZ}^{\mathrm{qf}}$ in parallel (see Fig. 3e), we realize a logical, transversal $\mathrm{CZ}^{\mathrm{qf}}$. This enables fault-tolerant qubit-fermion quantum computation. In particular, we will discuss in the next session that this has immense utility for implementing modern quantum algorithms for simulating fermions.
@@
Qubit only Fermion-qubit
fSWAP [14] FFFT [39] FFFT
Depth $\mathcal{O}(N)$ $\mathcal{O}(N\log{N})$ $\mathcal{O}(\log{N})$
Cliffords $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2}\log(N))$ $\mathcal{O}(N\log{N})$
Figure 4: Fermion-qubit logical quantum simulation primitives and comparison to the qubit-only approach. a, Qubit-controlled arbitrary-angle interaction gate. The box with label “$\theta/4$” is a qubit gate $\exp(-i(\theta^{\prime}/2)X)$ with $\theta^{\prime}=\theta/4$, synthesised from discrete gates using the Solovay-Kitaev algorithm. b, Qubit-controlled arbitrary-angle braiding gate. The gray box is the qubit-controlled braiding gate $\exp(-(\pi/4)Z\tilde{\gamma}_{i}\gamma_{j})$. c, FFFT circuit implemented using logical fermions. Crossing lines indicate swapping, implemented by movement of the atoms. d, Corresponding FFFT circuit implemented using logical qubits, c.f. Ref. [40]. e, Decomposition of the n-mode FFFT into an arbitrary-angle phase gate $\exp(-i\theta n)$ with $\theta=\theta_{k}^{n}=2\pi k/n$ and a two-mode FFFT. f, Compilation of the two-mode FFFT into fermionic fault-tolerant gates. g, Depth and Clifford complexity of qubit-only and fermion-qubit algorithms for simulating the kinetic energy term in crystalline materials, i.e. the first term in Eq. (7).
A prime area of application for fermionic quantum computing is simulating electrons in molecules and materials. In second quantized approaches, the continuous-space wavefunctions of electrons are cast into a discrete lattice model through a choice of basis, which leads to a Hamiltonian coupling fermionic operators as $H_{\mathrm{qchem}}=\sum_{ij}J_{ij}c^{\dagger}_{i}c_{j}+\sum_{ijkl}V_{ijkl}c^{\dagger}_{i}c^{\dagger}_{j}c_{k}c_{l}$, with $J_{ij}$ and $V_{ijkl}$ constants that depend on the choice of basis. This generally leads to a number of Hamiltonian terms scaling as $\sim N^{4}$ with the basis size $N$. For a Trotterized simulation, this implies the same scaling for the number of gates per Trotter step. However, for crystalline materials, the plane-wave-dual basis [12] enables writing the chemistry Hamiltonian as
@@
$\quad+\frac{1}{2}\sum_{ij}U_{ij}\left(n_{i}-\frac{1}{2}\right)\left(n_{j}-\frac{1}{2}\right),$ (7)
where $\epsilon_{i}$, $\mu_{i}$, and $U_{ij}$ are classically calculable coefficients and $\mathrm{FFFT}$ denotes the fermionic fast Fourier transform  [40]. This formulation hence reduces the scaling of the number of terms to $\sim N^{2}$. Efficient algorithms for dynamics, as well as for ground and finite-temperature state preparation, in addition require controlling this Hamiltonian on a qubit [10, 11].
Our next goal is therefore to implement qubit-controlled arbitrary-angle evolution under each of the terms in Eq. (7) individually. Time evolution under the Hamiltonian $\epsilon_{i}n_{i}$ until time $t$ is implemented by replacing the T gate in Fig. 3h with an approximate arbitrary-angle rotation of angle $\epsilon_{i}t$, where $t$ corresponds to time, synthesized from discrete gates with e.g. the Solovay-Kitaev algorithm. The interaction gate $\exp\left(-iZ\frac{\theta}{2}(n_{i}-\frac{1}{2})(n_{j}-\frac{1}{2})\right)$ is implemented by the circuit shown in Fig. 4a, which maps the arbitrary-angle single-qubit gate onto the fermions using $\mathrm{CZ}^{\mathrm{qf}}$  gates. A related gadget shown in Fig. 3b realizes an arbitrary-angle qubit-controlled braiding gate $\exp(-\frac{\theta}{2}Z\tilde{\gamma}_{i}\gamma_{j})$, where again $\mathrm{CZ}^{\mathrm{qf}}$  gates map the phase gate from the qubit to a fermion. The $\mathrm{BRAID}$s then “spread” the gate onto the second fermion. Two such gadgets are combined to yield the qubit-controlled hopping gate via $\exp\left(iJ_{ij}\Delta tZ\left(c_{i}^{\dagger}c_{j}+h.c.\right)\right)=\exp(-\frac{J\Delta t}{2}Z\tilde{\gamma}_{j}\gamma_{i})\exp(-\frac{J\Delta t}{2}Z\tilde{\gamma}_{i}\gamma_{j})$. These arbitrary-angle gates also enable time evolution under the more general Hamiltonian $H_{\mathrm{qchem}}$ by using the compilations in Ref. [27]. Finally, we need to implement the FFFT, which may be reduced to the 1D FFFT circuit shown in Fig. 4c, c.f. Ref. [40]. In analogy to the classical fast Fourier transform, it uses a divide-and-conquer reduction to two-site Fourier transforms $F^{k}_{N}$ (Fig. 4e, f) whose phases depend on the number of sites $N$ and an index $k$. Despite the fact that it is a 1D transformation, it requires non-local connectivity, which we can simply implement by moving the tweezers.

### v2

# Fault-tolerant fermionic quantum computing

@@
Fault-tolerant fermionic quantum computing
@@
(July 16, 2025)
Simulating the dynamics of electrons and other fermionic particles in quantum chemistry [1], materials science [2], and high-energy physics [3] is one of the most promising applications of fault-tolerant quantum computers. However, the overhead in mapping time evolution under fermionic Hamiltonians to qubit gates [4, 5, 6, 7] renders this endeavor challenging [8, 9]. We introduce fermion-qubit fault-tolerant quantum computing, a framework which removes this overhead altogether. Using native fermionic operations we first construct a repetition code which corrects phase errors only. Within a fermionic color code, which corrects for both phase and loss errors, we then realize a universal fermionic gate set, including transversal fermionic Clifford gates. Interfacing with qubit color codes we introduce qubit-fermion fault-tolerant computation, which allows for qubit-controlled fermionic time evolution, a crucial subroutine in state-of-the-art quantum algorithms for simulating fermions [10, 11, 12, 13]. As an application, we consider simulating crystalline materials [12], finding an exponential improvement in circuit depth for a single time step from $\mathcal{O}(N)$ to $\mathcal{O}(\log(N))$ with respect to lattice site number $N$ while retaining a fermion-site count of $\tilde{\mathcal{O}}(N)$, implying a linear-in-$N$ end-to-end gate depth for simulating materials, as opposed to quadratic in previous approaches. We also introduce a fermion-inspired qubit algorithm with $O(\mathrm{log}(N)$ depth, but which uses a prohibitive number of additional ancilla qubits compared to fermionic hardware. We show how our framework can be implemented in neutral atoms, overcoming the apparent inability of neutral atoms to implement non-number-conserving gates. Our work opens the door to fermion-qubit fault-tolerant quantum computation in platforms with native fermions such as neutral atoms, quantum dots and donors in silicon, with applications in quantum chemistry, material science, and high-energy physics.
@@
Solving the dynamics of many strongly-interacting fermionic quantum particles is central to understanding the dynamics of quarks in the early universe, the reaction dynamics of molecules, and high-temperature superconductivity. While qubit quantum computers promise to help solve these open problems, simulating the evolution of fermions encoded in qubits encounters a large overhead in both the number of two-qubit gates and the circuit depth [5, 14], which means that a solution of fermionic problems with near-term [8, 9] and even fault-tolerant quantum computers [15, 12] is extremely challenging. Quantum simulation experiments with cold fermionic atoms have made major progress in understanding continuum [16, 17, 18, 19, 20, 21] and lattice systems [22, 23, 24, 25]. These experiments have made steps towards digital computation by trapping individual fermionic atoms in optical tweezers [26, 27]. Additionally, qubits have been encoded into pairs of fermions [28]. Quantum simulations using trapped electrons in the solid state have also been demonstrated [29, 30, 31]. However, turning these fermionic platforms from a tool of scientific exploration into a computational tool requires error-correction and fault-tolerance techniques, for which no practical scheme has been found so far.
Here, we close this gap by introducing a framework for fermion-to-fermion fault-tolerant quantum computation. We first prove a theorem that fermionic error correction is not possible using only number-conserving gates, which are most easily implemented in experimental platforms. We then show how to avoid this theorem by using a non-number-conserving gate and concepts from Majorana error correction. This enables us to explicitly show how to perform a universal fermionic gate set including fault-tolerant transversal gates. To enable usage of state-of-the-art quantum algorithms requiring qubit-controlled gates, we introduce qubit-fermion fault-tolerant computation. We apply our framework to time evolution under the Hamiltonian of a crystalline material and study the advantages of our framework compared to all-qubit quantum computation. Finally, we show how to realize these codes and their operations in neutral atoms by removing the previously assumed restriction to number-conserving gates in neutral atoms and introducing the physics behind the implementation of a non-number-conserving braiding gate.
@@
The proof (see Methods) relies on the fact that codestates that are eigenstates of the total fermion number operator need to be eigenstates with the same number eigenvalue due to the Knill-Laflamme conditions. This means that logical operators cannot change the total fermion number and therefore need to be even weight in fermionic creation/annihilation operators. Hence, logical operators are bosonic.
@@
For implementing active error correction, we need to measure where errors happened and correct for them, which we show how to do in Fig. 2b. We use our qubit-fermion gate to map a measurement of the fermion density $n$ to a measurement of the phase of the qubit, which we transform to a measurement of the population of the qubit using Hadamard gates. To measure the stabilizer instead of $n$ on the fermions we conjugate the qubit-fermion gate with two braiding gates as well as phase gates. An ancilla qubit is used for each of the stabilizers. Dependent on the measurement outcome, $\mathrm{Z}^{\mathrm{f}}$  gates are performed on the fermions to correct for the errors. This scheme shows the first advantage of using qubits: destructive qubit readout is faster than a non-destructive $n$ measurement [40]. To prepare a logical codestate |$\ket{0}_{\mathrm{L}}$, we use a chain of braiding gates—the $\mathrm{S}^{\dagger}$ gates effectively braid the two Majoranas on the same physical site, Fig. 2c.
@@
We now show how to construct an error-correcting code correcting for both phase errors $E=n$ and atom-loss errors $E=p$ and show how to perform fault-tolerant gadgets. In the context of encoding qubits in Majorana nanowire platforms, Bravyi, Terhal, and Leemhuis [35] noted that a particular class of qubit error correcting codes, so-called weakly self-dual Calderbank-Shor-Steane (CSS) codes, can be converted into a Majorana stabilizer code by replacing qubit operators in the stabilizers by Majorana fermions. Motivated by the prospect of high error-resilience for codes with odd-weight logical operators, they noted that for codes with an odd number of physical Majorana fermions, a single odd-weight logical operator can be encoded. This operator is given by the total parity, i.e. the product of all Majorana operators. Our key observation is that, because there is an odd number of physical Majorana fermions, this logical operator is in fact a logical Majorana fermion operator. Even more importantly, while in Majorana nanowire platforms, a complex physical fermion would have to be encoded in two logical Majoranas which are encoded in spatially distinct physical Majoranas, our framework using complex physical fermions enables the encoding of two such Majorana codes, and therefore one complex logical fermion, in the same set of physical fermions. It turns out that a general prescription to constructing complex fermion codes from weakly self-dual CSS codes is to replace the $Z$ operators in the qubit code stabilizers by $\gamma$ operators and the qubit $X$ operators by $\tilde{\gamma}$ operators, see Fig. 3a. The logical Majorana operators are then given by the two parities $\gamma^{\mathrm{L}}=\prod_{i=1}^{N}\gamma_{i}$ and $\tilde{\gamma}^{\mathrm{L}}=\prod_{i=1}^{N}\tilde{\gamma}_{i}$, from which we construct a single logical fermion operator $c^{\dagger}=\frac{1}{2}\left(\gamma^{\mathrm{L}}+i\tilde{\gamma}^{\mathrm{L}}\right)$. As an example, we consider the family of triangular honeycomb Majorana color codes, which encodes a single logical fermion with odd distance $d$ into $N=(3d^{2}+1)/4$ physical fermions. In particular, $d=3$ is the fermionic version of the Steane code [41], Fig. 3a. Such codes can correct $(d-1)/2$ loss or phase errors.
@@
Figure 3: Fermion-qubit fault-tolerant quantum computation. a, Construction of the fermionic Steane code, the smallest member of the family of color codes we consider. Stabilizers are obtained by replacing products of $X$ ($Z$) operators in the qubit code with products of $\gamma$ ($\tilde{\gamma}$), respectively. Corners on areas with the same color are part of a stabilizer as shown for the bottom left area. b, Mid-circuit measurement of $-\gamma_{1}\gamma_{2}\gamma_{3}\gamma_{4}$ (counting sites from bottom to top). By inverting the $\mathrm{BRAID}$s (i.e. $\mathrm{BRAID}_{12}\rightarrow\mathrm{BRAID}_{21}$), we measure $-\tilde{\gamma}_{1}\tilde{\gamma}_{2}\tilde{\gamma}_{3}\tilde{\gamma}_{4}$, the other stabilizer. c, Neutral-atom implementation of transversal $\mathrm{CZ}^{\mathrm{qf}}$  gates, performed by interleaving a logical qubit and fermion and applying a global Rydberg laser. This implements the circuit shown in subfigure e. d, Transversal $\mathrm{CZ}^{\mathrm{f}}$. This is implemented by the logical qubit in subfigure c with a logical fermion. e, Transversal qubit-fermion $\mathrm{CZ}^{\mathrm{qf}}$. Transversal f, $\mathrm{BRAID}$  and g, $\mathrm{S}^{\mathrm{f}}$  gates are similarly implemented by parallel operations. h, $\mathrm{T}^{\mathrm{f}}$gate. Implemented by using a T gate applied on an ancilla qubit, which is then mapped onto the fermion. All the operations shown here are straightforwardly generalized to codes with larger (odd) $d$, see Methods.
@@
Of central utility is the fact that we can use qubits for controlled operations. One example is the measurement of stabilizers using the Hadamard test, as shown in Fig. 3b. From these stabilizers, we can reconstruct whether a $\tilde{\gamma}$ or $\gamma$ error (or both) happened on a fermion. Such single-Majorana errors are corrected by coupling an ancilla site $a$ prepared in |$\ket{0}$ to the site with the error, acting with $(\mathrm{BRAID}_{ai})^{2}=\gamma_{i}\tilde{\gamma}_{a}$ and then measuring the ancilla. The measurement circuit in Fig. 3b is made fault-tolerant by for example using the Shor scheme, i.e. replacing the single ancillary qubit with a set of qubits in a cat state and repeatedly measuring the stabilizers. This stabilizer measurement is also used for fault-tolerant preparation of the logical codestates from a trivial starting state. Note that the correction of stabilizers is optional for both state preparation and error correction—instead of enforcing stabilizers to be +1, their value can instead be tracked [42]. Moreover, we expect that error correction only needs to be applied after each gate, and not during movement, as demonstrated in neutral-atom qubit quantum computers [38].
@@
Magically, this fermion Steane code yields highly intuitive and easy-to-implement fault-tolerant logical gates. Transversal logical gates are implemented, similar to qubit Steane codes, by applying the corresponding physical gate in parallel on all atoms, c.f. Fig. 3c-g, see Methods. This set of gates turns out to be the Majorana Clifford gates $\mathrm{BRAID}$, $\mathrm{CZ}^{\mathrm{f}}$, $\mathrm{S}^{\mathrm{f}}$, which in analogy to qubit Clifford gates map a single Majorana operator string to another single Majorana string. These transversal Clifford gates are inherently fault-tolerant as an error on one physical codeblock can only spread to one physical fermion in the other codeblock.
@@
One of the key tools enabled by our approach is the interfacing of our fermionic Steane code with a qubit Steane code (with matching $N$) using the qubit-fermion $\mathrm{CZ}^{\mathrm{qf}}$ gate: by applying physical $\mathrm{CZ}^{\mathrm{qf}}$ in parallel (see Fig. 3e), we realize a logical, transversal $\mathrm{CZ}^{\mathrm{qf}}$. This enables fault-tolerant qubit-fermion quantum computation. In particular, we will discuss in the next section that this has immense utility for implementing modern quantum algorithms for simulating fermions.
@@
Qubits Fermions
fSWAP [44] pFFFT+pINT FFFT+pINT
Depth $\mathcal{O}(N)$ $\mathcal{O}(\log{N})$ $\mathcal{O}(\log{N})$
Ancillas 0 $\mathcal{O}(N^{2})$ + $\mathcal{O}(N\log{N})$ 0 + $\mathcal{O}(N\log{N})$
Volume $\mathcal{O}(N^{2})$ $\tilde{\mathcal{O}}(N^{2})$ $\tilde{\mathcal{O}}(N)$
Figure 4: Fermion-qubit logical quantum simulation primitives and comparison to the qubit-only approach. a, Qubit-controlled arbitrary-angle interaction gate. The box with label “$\theta/4$” is a qubit gate $\exp(-i(\theta^{\prime}/2)X)$ with $\theta^{\prime}=\theta/4$, synthesised from discrete gates using the Solovay-Kitaev algorithm. See Fig. 1 for our convention of the labeling of the braiding gate. d, Qubit-controlled arbitrary-angle braiding gate. The gray box is the qubit-controlled braiding gate $\exp(-(\pi/4)Z\tilde{\gamma}_{i}\gamma_{j})$. c, FFFT circuit implemented using logical fermions. Crossing lines indicate swapping, implemented by movement of the atoms. d, Corresponding FFFT circuit implemented using logical qubits, c.f. Ref. [45]. e, Decomposition of the n-mode FFFT into an arbitrary-angle phase gate $\exp(-i\theta n)$ with $\theta=\theta_{k}^{n}=2\pi k/n$ and a two-mode FFFT. f, Compilation of the two-mode FFFT into fermionic fault-tolerant gates. g, Scaling of depth, number of ancillas and circuit volume (defined as the product of depth and the number of ancillas) with the number of sites $N$ for a single Trotter step of the plane-wave basis Hamiltonian in Eq. (7). We compare the lowest-depth previous method [12], our qubit-only and our fermion-qubit algorithms. “pINT” refers to a parallelized interactions scheme using the quantum FFT [46], “pFFFT” to our version of the FFFT which using the parallelization of CZ circuits introduced in Ref. [47].
A prime area of application for fermionic quantum computing is simulating electrons in molecules and materials. In second quantized approaches, the continuous-space wavefunctions of electrons are cast into a discrete lattice model through a choice of basis, which leads to a Hamiltonian coupling fermionic operators as $H_{\mathrm{qchem}}=\sum_{ij}J_{ij}c^{\dagger}_{i}c_{j}+\sum_{ijkl}V_{ijkl}c^{\dagger}_{i}c^{\dagger}_{j}c_{k}c_{l}$, with $J_{ij}$ and $V_{ijkl}$ constants that depend on the choice of basis. This generally leads to a number of Hamiltonian terms scaling as $\sim N^{4}$ with the basis size $N$. For a Trotterized simulation, this implies the same scaling for the number of gates per Trotter step. However, the plane-wave-dual basis [12], which is particularly suitable for simulating crystalline materials, enables writing the chemistry Hamiltonian as
@@
$\quad+\sum_{i<j}U_{ij}\left(n_{i}-\frac{1}{2}\right)\left(n_{j}-\frac{1}{2}\right),$ (7)
where $\epsilon_{i}$, $\mu_{i}$, and $U_{ij}$ are classically calculable coefficients and $\mathrm{FFFT}$ denotes the fermionic fast Fourier transform  [45]. This formulation reduces the scaling of the number of terms to $\sim N^{2}$. Efficient algorithms for dynamics, as well as for ground and finite-temperature state preparation, in addition often require controlling this Hamiltonian on a qubit [10, 11].
Our next goal is therefore to implement qubit-controlled arbitrary-angle evolution under each of the terms in Eq. (7) individually. Time evolution under the Hamiltonian $\epsilon_{i}n_{i}$ until time $t$ is implemented by replacing the T gate in Fig. 3h with an approximate arbitrary-angle rotation of angle $\epsilon_{i}t$, where $t$ corresponds to time, synthesized from discrete gates with e.g. the Solovay-Kitaev algorithm. The interaction gate $\exp\left(-iZ\frac{\theta}{2}(n_{i}-\frac{1}{2})(n_{j}-\frac{1}{2})\right)$ is implemented by the circuit shown in Fig. 4a, which maps the arbitrary-angle single-qubit gate onto the fermions using $\mathrm{CZ}^{\mathrm{qf}}$  gates. A related gadget shown in Fig. 3b realizes an arbitrary-angle qubit-controlled braiding gate $\exp(-\frac{\theta}{2}Z\tilde{\gamma}_{i}\gamma_{j})$, where again $\mathrm{CZ}^{\mathrm{qf}}$  gates map the phase gate from the qubit to a fermion. The $\mathrm{BRAID}$s then “spread” the gate onto the second fermion. Two such gadgets are combined to yield the qubit-controlled hopping gate via $\exp\left(iJ_{ij}\Delta tZ\left(c_{i}^{\dagger}c_{j}+h.c.\right)\right)=\exp(-\frac{J\Delta t}{2}Z\tilde{\gamma}_{j}\gamma_{i})\exp(-\frac{J\Delta t}{2}Z\tilde{\gamma}_{i}\gamma_{j})$. These arbitrary-angle gates also enable time evolution under the more general Hamiltonian $H_{\mathrm{qchem}}$ by using the compilations in Ref. [36]. Finally, we need to implement the FFFT, which may be reduced to the 1D FFFT circuit shown in Fig. 4c, c.f. Ref. [45]. In analogy to the classical fast Fourier transform, the FFFT uses a divide-and-conquer reduction to two-site Fourier transforms $F^{k}_{N}$ (Fig. 4e, f) whose phases depend on the number of sites $N$ and an index $k$. Despite the fact that it is a 1D transformation, it requires non-local connectivity, which is implemented in neutral atoms by moving the tweezers [48].
@@
We now show that our approach yields a large advantage over currently known qubit algorithms for simulating Trotterized time evolution under the Hamiltonian of crystalline materials.